# Symplectic geometry of electrical networks

### v2

@@ Abstract.
In this paper we relate a well-known in symplectic geometry compactification of the space of symmetric bilinear forms on the Lagrangian Grassmannian to the specific compactifications of the space of electrical networks in the disc obtained in [11], [4] and [3]. In particular, we state an explicit connection between these works and describe some of the combinatorics developed there in the language of symplectic geometry. We also show that the combinatorics of the concordance vectors forces the uniqueness of the symplectic form, such that corresponding points of the Grassmannian are isotropic. We define a notion of Lagrangian concordance which provides a construction of the compactification of the space of electrical networks in the positive part of the Lagrangian Grassmannian bypassing the construction from [11].
@@ 1.2. Acknowledgments
Research of B. B. was partially supported by the ISF grant 876/20. For V. G. this article is an output of a research project implemented as part of the Basic Research Program at the National Research University Higher School of Economics (HSE University). Working on this project V. G. also visited the Max Plank Institute of Mathematics in Bonn, Germany in the summer 2023 and BIMSA in Beijing, China in the summer 2024. L. G. was partially supported by the Israel Science Foundation grant ISF-2480/20. Research of A. K. and B. B. on Sections 3 and 4 was supported by the Russian Science Foundation project No. 20-71-10110 (https://rscf.ru/en/project/23-71-50012) which finances the work of A.K. at P. G. Demidov Yaroslavl State University. Research of A. K. on Section 5 was supported by the state assignment of MIPT (project FSMG-2023-0013). We are greateful to Boris Feigin for useful discussions.
@@ 2.1. Electrical networks
We start with a brief introduction into the theory of electrical networks following [5] and [7]. An electrical network is a finite weighted undirected graph $\Gamma$ embedded into a disk, where the vertex set is divided into the set of boundary vertices and the set of interior vertices. The weight $w(e)$ of an edge is to be thought of a conductance of the corresponding edge. We will think of electrical current as a flow along edges, and each vertex has a voltage. These physical objects should satisfy two laws. Kirchhoff’s law says that for any interior vertex the total current through all adjacent edges equals zero. Ohm’s law says that if $e=(u,v)$ is an edge then we have
@@
where $w(e)$ is the conductance of the edge, $U(u),\ U(v)$ are the voltages of the vertices $u,\ v$ respectively and $I$ is the current from $u$ to $v$. We think of boundary vertices as of nodes to which a voltage is applied. It follows from the maximum principle for the discrete harmonic functions [6, Chapter 3.2] that fixed voltages on the boundary vertices give rise to the unique extension on all vertices. Voltages on the vertices lead to the electrical current flowing along the edges. From Kirchhoff’s law follows that to determine electrical current on any edge of the graph $\Gamma$, it is enough to know electrical currents outgoing from the boundary vertices. Therefore to completely determine physical properties of an electrical network it is enough to know the map sending a vector $v\in\mathbb{R}^{n}$ of boundary voltages to a vector of boundary currents. This map is linear and its matrix is called the response matrix $M_{R}$ which can be determined in terms of the network as it is proved in [5, Theorem 3.2].
We denote an electrical network by $e(\Gamma,\omega)$ or simply $e$, where $\Gamma$ is an underlying graph and $\omega:\text{vertex set of }\Gamma\to\mathbb{R}_{>0}$ is a weight (conductance) function. We denote by $E_{n}$ the set of electrical networks on $n$ boundary vertices.
@@ 2.2. Compactification of the spaces of bilinear forms and electrical networks
The set of all Lagrangian subspaces of $W_{n}$ is the set of points of an algebraic variety $LG(n)$, called the Lagrangian Grassmannian. Let $H$ be a hyperspace in $W_{n}$ and $H^{\perp}$ be its orthogonal with respect to the symplectic form $\Lambda$. It is a standard exercise in the linear symplectic geometry to show that $\mathrm{dim}\,H^{\perp}=1$, $H^{\perp}\subset H$ and if $L$ is a Lagrangian subspace in $W_{n}$ then $H^{\perp}\subset L$ and $\dim L\cap H=n-1.$ Also $H^{\perp}\cap(L\cap H)\neq\emptyset$, therefore
@@ 2.3. Positive compactifications of electrical networks
Recall that we denote by $E_{n}$ the set of electrical networks with $n$ boundary vertices. We denote by $\overline{E}_{n}$ the corresponding set of cactus networks.
@@
The totally nonnegative Grassmannian is the locus in the Grassmannian with the Plucker coordinates with the same sign.
@@
It was proven by Chepuri-George-Speyer [4, Theorem 1.5] and Bychkov-Gorbounov-Kazakov-Talalaev [3, Theorem 3.3] that in fact the image of $\overline{E}_{n}$ under the Lam’s embedding lies in the totally nonnegative Lagrangian Grassmannian. Now we are going to describe the embedding $\overline{E}_{n}$ into the totally nonnegative Grassmannian $Gr(n-1,2n)_{\geq 0}$ following [3]. This method allows to refine Theorem 2.7 since it proves that the image of $\overline{E}_{n}$ lies in the Lagrangian Grassmannian $LG(n-1,V)$ for a certain $22n-$ 2-dimensional subspace $V$. We will describe the second approach from [4] in Section 3.
@@
[3, Theorem 3.3] The row space of $\Omega_{n}(e)$ defines the same point in the $Gr_{\geq 0}(n-1,2n)$ as the point defined by $e(\Gamma,\omega)$ under the Lam’s embedding. More precisely, the set of maximal minors of the matrix $\Omega_{n}(e)$ is related to the choice of Plucker coordinates $\Delta_{I}^{\bullet}$ in Lam’s embedding as $\Delta_{I}(\Omega_{n}(e))=\frac{\Delta_{I}^{\bullet}}{L_{unc}}$.
@@ 4.3. Lagrangian concordance
Figure 6. Non-crossing partition $\sigma=(\bar{1}|\bar{2},\bar{5},\bar{8}|\bar{3}|\bar{4}|\bar{6},\bar{7})$ of $[\bar{n}]$, $\widetilde{\sigma}$ of $[\widetilde{n}]$ and $(\sigma|\widetilde{\sigma})$.
@@
Let $\sigma$ be a non-crossing partition and $\overline{B}_{j}=\{(\overline{i}_{1},\overline{i}_{2},\dots,\overline{i}_{k})\in\sigma|\ k>1\}$ be its connected component, then extension $\mathrm{ext}(\overline{B}_{k})$ is the following sequence of indices:
@@
where $\mathrm{ext}(\overline{i}_{l}\overline{i}_{l-1})=(2i_{l}-1,2i_{l}+1,2i_{l}+3,\dots,2i_{l}-2).$
@@
where $\mathrm{ext}(\widetilde{i}_{l}\widetilde{i}_{l-1})=(2i_{l},2i_{l}+2,2i_{l}+4,\dots,2i_{l}-2).$
@@
Let $\sigma$ be a non-crossing partition and $\overline{B}_{j_{1}}\in\sigma$ and $\overline{B}_{j_{2}}\in\sigma$ be two its connected components, then we will say that $\overline{B}_{j_{2}}$ envelopes $\overline{B}_{j_{1}}$ if there is $(\overline{i}_{l-1}\overline{i}_{l})\in\overline{B}_{j_{2}}$ such that $\overline{B}_{j_{1}}\subset\mathrm{ext}(\overline{i}_{l-1}\overline{i}_{l})$. For $\overline{B}_{j_{2}}$ envelopes $\overline{B}_{j_{1}}$ we will use the notation $\overline{B}_{j_{1}}\in\mathrm{env}(\overline{B}_{j_{2}})$.
@@
We will say that a subset $I\subset[2n-2],\ I=|n-1|$ is Lagrangian concordant with $\sigma$ if $I$ shares exactly one element with each subcomponent of $\mathrm{Lext}(\sigma).$
@@ 5.1. Parametrizations via effective resistance matrices
aaa $b$$c$$a$$\bar{2}$$\bar{1}$$\bar{3}$$\Gamma$aaa $\frac{1}{b}$$\frac{1}{c}$$\frac{1}{a}$$\bar{2}$$\bar{1}$$\bar{3}$$\Gamma^{*}$
Figure 9.
@@
(21) $\Omega_{n,R}(e)=\left(\begin{matrix}1&-m_{12}&0&m_{13}&0&-m_{14}&0&\ldots&(-1)^{n}&(-1)^{n}m_{11}\\ 1&m_{22}&1&-m_{23}&0&m_{24}&0&\ldots&0&(-1)^{n+1}m_{12}\\ 0&-m_{23}&1&m_{33}&1&-m_{34}&0&\ldots&0&(-1)^{n+2}m_{13}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\end{matrix}\right),$
where $m_{ij}=\frac{1}{2}(R_{i,j}+R_{i+1,j+1}-R_{i,j+1}-R_{i+1,j})$.
@@
Consider $e\in E_{n}$ and its dual network $e^{*}$. By Lemma 5.3 $\Omega_{n}(e^{*})s^{-1}$ and $\Omega_{n}(e)$ define the same point of the Grassmannian $Gr_{\geq 0}(n-1,2n)$.
@@
$\Omega_{n}(e^{*})s^{-1}=\left(\begin{matrix}1&-x^{*}_{12}&0&x^{*}_{13}&0&-x^{*}_{14}&0&\ldots&(-1)^{n}&(-1)^{n}x^{*}_{11}\\ 1&x^{*}_{22}&1&-x^{*}_{23}&0&x^{*}_{24}&0&\ldots&0&(-1)^{n+1}x^{*}_{12}\\ 0&-x^{*}_{23}&1&x^{*}_{33}&1&-x^{*}_{34}&0&\ldots&0&(-1)^{n+2}x^{*}_{13}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\end{matrix}\right),$
@@
To finish the proof it remains to notice that $\Delta_{135\dots 2n-3}(\Omega^{\prime}_{n}(e^{*})s^{-1})=1$ and $\Delta_{135\dots 2n-3}^{\bullet}=L_{123\dots n}$ (see Example 2.8). ∎
@@
Now, consider the dual to $P(\sigma)$ hollow cactus network $P(\widetilde{\sigma})\in\overline{E}_{n}$ with the boundary defined by the non-crossing partition $\widetilde{\sigma}=({\widetilde{B}_{1},\widetilde{B}_{2},\dots,\widetilde{B}_{l}})$ dual to $\sigma$. Then let us repeat the above arguments for each connected component $\widetilde{B}_{i}\in\widetilde{\sigma}:\ |\widetilde{B}_{i}|>1$, then we have: $\bigoplus\limits_{\widetilde{B}_{i}\in\widetilde{\sigma}:\ |\widetilde{B}_{i}|>1}V(\widetilde{B}_{i})\subset p(\widetilde{\sigma}).$ Note that by Remark 5.2 we have $P(\widetilde{\sigma})=(P(\sigma))^{*}$. Let us denote by $(p(\sigma))^{*}$ the point of $Gr_{\geq 0}(n-1,2n)$ corresponding to $(P(\sigma))^{*}$. Thus we have $\bigoplus\limits_{\widetilde{B}_{i}\in\widetilde{\sigma}:\ |\widetilde{B}_{i}|>1}V(\widetilde{B}_{i})\subset(p(\sigma))^{*}.$ Identify the space $V(\widetilde{B}_{i})$ with a matrix representing it: it is a matrix whose rows define analogously to vectors $v_{\bar{j_{l}}\bar{j}_{l+1}}$ with the substitution $\bar{i}$ to $\widetilde{i}$ in the description below (28). Then we obtain: $\bigoplus\limits_{\widetilde{B}_{i}\in\widetilde{\sigma}:\ |\widetilde{B}_{i}|>1}V(\widetilde{B}_{i})s^{-1}\subset(p(\sigma))^{*}s^{-1}.$ Then by Lemma 5.3 we have: $\bigoplus\limits_{\widetilde{B}_{i}\in\widetilde{\sigma}:\ |\widetilde{B}_{i}|>1}V(\widetilde{B}_{i})s^{-1}\subset p(\sigma).$
And also for any $\widetilde{B}_{i}=\{\widetilde{j}_{1},\widetilde{j}_{2},\dots,\widetilde{j}_{k_{i}}\},\ |\widetilde{B}_{i}|=k_{i}>1$ we have that $V(\widetilde{B}_{i})s^{-1}=\mathrm{span}(v_{\widetilde{j}_{1}\widetilde{j}_{2}},v_{\widetilde{j}_{2}\widetilde{j}_{3}},\dots,v_{\widetilde{j}_{k_{i}-1}\widetilde{j}_{k_{i}}}),$ where

### v4

@@ Abstract.
In this paper we relate a well-known in symplectic geometry compactification of the space of symmetric bilinear forms considered as a chart of the Lagrangian Grassmannian to the specific compactifications of the space of electrical networks in the disc obtained in [11], [4] and [3]. In particular, we state an explicit connection between these works and describe some of the combinatorics developed there in the language of symplectic geometry. We also show that the combinatorics of the concordance vectors forces the uniqueness of the symplectic form, such that corresponding points of the Grassmannian are isotropic. We define a notion of Lagrangian concordance which provides a construction of the compactification of the space of electrical networks in the positive part of the Lagrangian Grassmannian bypassing the construction from [11].
@@ 1.2. Acknowledgments
Research of B. B. was partially supported by the ISF grant 876/20. For V. G. this article is an output of a research project implemented as part of the Basic Research Program at the National Research University Higher School of Economics (HSE University). Working on this project V. G. also visited the Max Plank Institute of Mathematics in Bonn, Germany in the summer 2023 and BIMSA in Beijing, China in the summer 2024. L. G. was partially supported by the Israel Science Foundation grant ISF-2480/20. Research of A. K. and B. B. on Sections 3 and 4 was supported by the Russian Science Foundation project No. 20-71-10110 (https://rscf.ru/en/project/23-71-50012) which finances the work of A.K. at P. G. Demidov Yaroslavl State University. Research of A. K. on Section 5 was supported by the state assignment of MIPT (project FSMG-2023-0013). We are greateful to Boris Feigin for useful discussions. We also thank the anonymous referee for valuable comments.
@@ 2.1. Electrical networks
We start with a brief introduction into the theory of electrical networks following [5] and [7]. An electrical network is a finite weighted undirected graph $\Gamma$ embedded into a disk, where the vertex set is divided into the set of clockwise enumerated boundary vertices and the set of interior vertices. The weight $w(e)$ of an edge is to be thought of a conductance of the corresponding edge. We will think of electrical current as a flow along edges, and each vertex has a voltage. These physical objects should satisfy two laws. Kirchhoff’s law says that for any interior vertex the total current through all adjacent edges equals zero. Ohm’s law says that if $e=(u,v)$ is an edge then we have
@@
where $w(e)$ is the conductance of the edge, $U(u),\ U(v)$ are the voltages of the vertices $u,\ v$ respectively and $I$ is the current from $u$ to $v$. We think of boundary vertices as of nodes to which a voltage is applied. It follows from the maximum principle for the discrete harmonic functions [6, Chapter 3.2] that fixed voltages on the boundary vertices give rise to the unique extension on all vertices. Voltages on the vertices lead to the electrical current flowing along the edges. From Kirchhoff’s law follows that to determine electrical current on any edge of the graph $\Gamma$, it is enough to know electrical currents outgoing from the boundary vertices. Therefore to completely determine physical properties of an electrical network it is enough to know the map sending a vector $v\in\mathbb{R}^{n}$ of boundary voltages to a vector of boundary currents. This map is linear and its matrix is called the response matrix $M_{R}$ which can be determined in terms of the network as it is proved in [5, Theorem 3.2]. Notice that two electrical networks share the same response matrix if and only if they can be transformed to each other with electrical transformations , see [11, Section 2.4].
We denote an electrical network by $e(\Gamma,\omega)$ or simply $e$, where $\Gamma$ is an underlying graph and $\omega:\text{vertex set of }\Gamma\to\mathbb{R}_{>0}$ is a weight (conductance) function. We denote by $E_{n}$ the set of electrical networks on $n$ boundary vertices up to the electrical transformations.
@@ 2.2. Compactification of the spaces of bilinear forms and electrical networks
The set of all Lagrangian subspaces of $W_{n}$ is the set of points of an algebraic variety $LG(n)$, called the Lagrangian Grassmannian. Let $H$ be a hyperspace in $W_{n}$ and $H^{\perp}$ be its orthogonal with respect to the symplectic form $\Lambda$. It is a standard exercise in the linear symplectic geometry to show that $\mathrm{dim}\,H^{\perp}=1$, $H^{\perp}\subset H$. If $L$ is a Lagrangian subspace in $W_{n}$ then $H^{\perp}$ is either a subspace in $L$ or $\mathrm{dim}\,H^{\perp}\cap L=0.$ In the first case $L$ must be a subspace of $H$ since the form $\Lambda$ is non degenerate, in the second case $\mathrm{dim}\,L\cap H=n-1$ since for non degenerate the maximal dimension of an isotropic subspace is $n$. Therefore in both these cases
@@ 2.3. Positive compactifications of electrical networks
A non-crossing partition $\sigma$ is a set partition of $[\bar{n}]:=\{\bar{1},\bar{2},\ldots,\bar{n}\}$ such that if $[\bar{n}]$ is arranged on a circle the convex hulls of its blocks are pairwise disjoint.
Equivalently, $\sigma$ is non-crossing if there is no $i<j<k<l$ with $ik$ and $jl$ belonging to different components of $\sigma$.
Non-crossing partitions can also be defined given a graph embedded into a disk.
Definition 2.2.
@@
Recall that we denote by $E_{n}$ the set of electrical networks with $n$ boundary vertices up to the electrical transformations, see [11, Section 2.4]. We denote by $\overline{E}_{n}$ the corresponding set of cactus networks up to the electrical transformations.
@@
The totally nonnegative Grassmannian is the locus in the Grassmannian with the Plucker coordinates of the same sign.
@@
Here we identify $Gr(n-1,2n)_{\geq 0}$ and its Plucker embedding. It was proven by Chepuri-George-Speyer [4, Theorem 1.5] and Bychkov-Gorbounov-Kazakov-Talalaev [3, Theorem 3.3] that in fact the image of $\overline{E}_{n}$ under the Lam’s embedding lies in the totally nonnegative Lagrangian Grassmannian. Now we are going to describe the embedding $\overline{E}_{n}$ into the totally nonnegative Grassmannian $Gr(n-1,2n)_{\geq 0}$ following [3]. This method allows to refine Theorem 2.8 since it proves that the image of $\overline{E}_{n}$ lies in the Lagrangian Grassmannian $LG(n-1,V)$ for a certain $22n-$ 2-dimensional subspace $V$. We will describe the second approach from [4] in Section 3.
@@
[3, Theorem 3.3] The row space of $\Omega_{n}(e)$ defines the same point in the $Gr_{\geq 0}(n-1,2n)$ as the point defined by $e(\Gamma,\omega)$ under the Lam’s embedding. More precisely, the set of maximal minors of the matrix $\Omega^{\prime}_{n}(e)$ obtained from $\Omega_{n}(e)$ by deleting the last row is related to the choice of Plucker coordinates $\Delta_{I}^{\bullet}$ in Lam’s embedding as $\Delta_{I}(\Omega_{n}(e))=\frac{\Delta_{I}^{\bullet}}{L_{unc}}$.
@@ 4.3. Lagrangian concordance
Figure 6. Non-crossing partition $\sigma=(\bar{1}|\bar{2},\bar{5},\bar{8}|\bar{3}|\bar{4}|\bar{6},\bar{7})$ of $[\bar{n}]$, $\widetilde{\sigma}$ of $[\widetilde{n}]$ and $(\sigma|\widetilde{\sigma})$ of $[2n]$.
@@
Let $\sigma$ be a non-crossing partition and $\overline{B}_{j}=\{(\overline{i}_{1},\overline{i}_{2},\dots,\overline{i}_{k})\in\sigma|\ k>1\}$ be its connected component, then extension $\mathrm{ext}(\overline{B}_{j})$ is the following sequence of indices:
@@
where $\mathrm{ext}(\overline{i}_{l}\overline{i}_{l+1})=(2i_{l}-1,2i_{l}+1,2i_{l}+3,\dots,2i_{l+1}-3).$
@@
where $\mathrm{ext}(\widetilde{i}_{l}\widetilde{i}_{l+1})=(2i_{l},2i_{l}+2,2i_{l}+4,\dots,2i_{l+1}-2).$
@@
Let $\sigma$ be a non-crossing partition and $\overline{B}_{j_{1}}\in\sigma$ and $\overline{B}_{j_{2}}\in\sigma$ be two its connected components of size greater than $1$, then we will say that $\overline{B}_{j_{2}}$ envelopes $\overline{B}_{j_{1}}$ if there is $(\overline{i}_{l-1}\overline{i}_{l})\in\overline{B}_{j_{2}}$ such that $\overline{B}_{j_{1}}\subset\mathrm{ext}(\overline{i}_{l-1}\overline{i}_{l})$. For $\overline{B}_{j_{2}}$ envelopes $\overline{B}_{j_{1}}$ we will use the notation $\overline{B}_{j_{1}}\in\mathrm{env}(\overline{B}_{j_{2}})$.
@@
We will say that a subset $I\subset[2n-2],\ I=|n-1|$ is Lagrangian concordant with $\sigma$ if $I$ shares exactly one element with each sub-component of $\mathrm{Lext}(\sigma).$
@@ 5.1. Parametrizations via effective resistance matrices
aaa $b$$\bar{2}$$\bar{1}$$\bar{3}$$c$$a$$\Gamma$aaa $\bar{2}$$\bar{1}$$\bar{3}$$\frac{1}{b}$$\frac{1}{a}$$\frac{1}{c}$$\Gamma^{*}$
Figure 9. Action of the operator $s^{-1}$ on the network
@@
(21) $\Omega_{n,R}(e)=\left(\begin{matrix}1&m_{11}&1&-m_{12}&0&m_{13}&0&\ldots\\ 0&-m_{21}&1&m_{22}&1&-m_{23}&0&\ldots\\ 0&m_{31}&0&-m_{32}&1&m_{33}&1&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{matrix}\right),$
where $m_{ij}=-\frac{1}{2}(R_{i,j}+R_{i+1,j+1}-R_{i,j+1}-R_{i+1,j})$.
@@
Consider $e\in E_{n}$ and its dual network $e^{*}$. By Lemma 5.3 $\Omega_{n}(e^{*})s$ and $\Omega_{n}(e)$ define the same point of the Grassmannian $Gr_{\geq 0}(n-1,2n)$.
@@
$\Omega_{n}(e^{*})s=\left(\begin{matrix}1&x^{*}_{11}&1&-x^{*}_{12}&0&x^{*}_{13}&0&\ldots\\ 0&-x^{*}_{21}&1&x^{*}_{22}&1&-x^{*}_{23}&0&\ldots\\ 0&x^{*}_{31}&0&-x^{*}_{32}&1&x^{*}_{33}&1&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{matrix}\right),$
@@
To finish the proof it remains to notice that $\Delta_{135\dots 2n-3}(\Omega^{\prime}_{n}(e^{*})s)=1$ and $\Delta_{135\dots 2n-3}^{\bullet}=L_{123\dots n}$ (see Example 2.9). ∎
@@
Now, consider the dual to $P(\sigma)$ hollow cactus network $P(\widetilde{\sigma})\in\overline{E}_{n}$ with the boundary defined by the non-crossing partition $\widetilde{\sigma}=({\widetilde{B}_{1},\widetilde{B}_{2},\dots,\widetilde{B}_{l}})$ dual to $\sigma$. Then let us repeat the above arguments for each connected component $\widetilde{B}_{i}\in\widetilde{\sigma}:\ |\widetilde{B}_{i}|>1$, then we have: $\bigoplus\limits_{\widetilde{B}_{i}\in\widetilde{\sigma}:\ |\widetilde{B}_{i}|>1}V(\widetilde{B}_{i})\subset p(\widetilde{\sigma}).$ Note that by Remark 5.2 we have $P(\widetilde{\sigma})=(P(\sigma))^{*}$. Let us denote by $(p(\sigma))^{*}$ the point of $Gr_{\geq 0}(n-1,2n)$ corresponding to $(P(\sigma))^{*}$. Thus we have $\bigoplus\limits_{\widetilde{B}_{i}\in\widetilde{\sigma}:\ |\widetilde{B}_{i}|>1}V(\widetilde{B}_{i})\subset(p(\sigma))^{*}.$ Identify the space $V(\widetilde{B}_{i})$ with a matrix representing it: it is a matrix whose rows define analogously to vectors $v_{\bar{j_{l}}\bar{j}_{l+1}}$ with the substitution $\bar{i}$ to $\widetilde{i}$ in the description below (28). Then we obtain: $\bigoplus\limits_{\widetilde{B}_{i}\in\widetilde{\sigma}:\ |\widetilde{B}_{i}|>1}V(\widetilde{B}_{i})s\subset(p(\sigma))^{*}s.$ Then by Lemma 5.3 we have: $\bigoplus\limits_{\widetilde{B}_{i}\in\widetilde{\sigma}:\ |\widetilde{B}_{i}|>1}V(\widetilde{B}_{i})s\subset p(\sigma).$
And also for any $\widetilde{B}_{i}=\{\widetilde{j}_{1},\widetilde{j}_{2},\dots,\widetilde{j}_{k_{i}}\},\ |\widetilde{B}_{i}|=k_{i}>1$ we have that $V(\widetilde{B}_{i})s=\mathrm{span}(v_{\widetilde{j}_{1}\widetilde{j}_{2}},v_{\widetilde{j}_{2}\widetilde{j}_{3}},\dots,v_{\widetilde{j}_{k_{i}-1}\widetilde{j}_{k_{i}}}),$ where